# MambaMIR: An Arbitrary-Masked Mamba for Joint Medical Image Reconstruction and Uncertainty Estimation

### v1

@@ 1 Introduction
Magnetic Resonance Imaging (MRI) and X-ray Computed Tomography (CT) are two representative medical imaging tools. Magnetic resonance imaging serves as a crucial non-invasive diagnostic tool, offering high-resolution and reproducible assessments of both structural and functional information, without exposure to radiation. Fast MRI is widely utilized to produce MR images from sub-Nyquist sampled k-space measurements, aiming to speed up the inherently slow data acquisition process [18, 10, 13]. This accelerated acquisition process can effectively reduce artifacts from voluntary and involuntary physiological movements during scanning. X-ray Computed Tomography is renowned for its ability to generate high-quality and detailed images. However, it also raises significant concerns regarding potential cancer risks due to radiation exposure [30]. Sparse-view CT (SVCT), which reconstructs images from far fewer projection views, can effectively reduce the radiation dose. Moreover, in applications such as C-arm CT or dental CT, where scanning time is predominantly constrained by the slower flat-panel detector rather than the gantry’s mechanical speed, sparse-view CT emerges as a promising solution to expedite the scanning process [24]. However, reducing the number of projection views during scanning can introduce severe artifacts in reconstructed images, which can significantly undermine clinical diagnosis. Therefore, for fast MRI, SVCT, and other medical image reconstruction tasks, a key research topic and challenge is to develop effective, efficient, and reliable reconstruction models. Therefore, for fast MRI, SVCT as well as other medical image reconstruction task, a key research topic and challenge is to develop effective, efficient and reliable reconstruction model.
The rapid advancement of artificial intelligence has propelled the development and widespread application of deep learning-based medical image reconstruction. Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) [3] represent two mainstream models that have achieved remarkable success in various computer vision tasks and are widely utilised in the medical imaging field.
@@
CNNs are effective feature extractors, taking the advantage of their hierarchical architecture and inductive bias, especially excelling at identifying local patterns. The shared weight mechanism renders them more parameter-efficient than multilayer perceptrons (MLPs). Various CNN-based medical image reconstruction models have been proposed for different modalities, including MRI [28, 1], CT [16, 9], Ultrasound [37], and Positron Emission Tomography (PET) [6]. However, CNNs typically exhibit local sensitivity and a lack of long-range dependency, which limits their ability to contextualise global features.
Vision Transformers [3], characterised by their large receptive fields and global sensitivity, often outperform CNNs in capturing extensive contextual information. Nonetheless, their significant computational demand, driven by the self-attention mechanism’s quadratic complexity with respect to sequence length, limits their practicality for medical image reconstruction, where images often have high resolution. Moreover, utilising larger patch sizes may compromise the reconstruction of detailed information [19]. Recent Transformer-based models for medical image reconstruction have sought to mitigate these limitations by: 1) Adopting a trade-off strategy that applies the self-attention mechanism within shifting windows rather than across the entire feature map [12]; 2) Constructing hybrid models that incorporate CNNs [2] or Swin Transformers [21], applying ViT blocks only within deep, low-resolution latent spaces [2, 15].
The rapid advancement of artificial intelligence has propelled the development and widespread application of deep learning (DL)-based medical image reconstruction. Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) [3] represent two mainstream models that have achieved remarkable success in various computer vision tasks and are widely utilized in the medical imaging field. CNNs, with their hierarchical architecture and inductive biases, excel at extracting features, especially in identifying local patterns. The shared weight mechanism renders them more parameter-efficient than multilayer perceptrons (MLPs). Various CNN-based medical image reconstruction models have been proposed for different modalities, including MRI [28, 1], CT [16, 9], Ultrasound [37], and Positron Emission Tomography (PET) [6]. However, CNNs typically exhibit local sensitivity and a lack of long-range dependency, which limits their ability to contextualize global features. Vision Transformers [3], characterized by their large receptive fields and global sensitivity, often outperform CNNs in capturing extensive contextual information. Nonetheless, their significant computational demand, driven by the self-attention mechanism’s quadratic complexity with respect to sequence length, limits their practicality for medical image reconstruction, where images often have high resolution. Moreover, utilizing larger patch sizes may compromise the reconstruction of detailed information [19]. Recent Transformer-based models for medical image reconstruction have sought to mitigate these limitations by: 1) Adopting a trade-off strategy that applies the self-attention mechanism within shifting windows rather than across the entire feature map [12]; 2) Constructing hybrid models that incorporate CNNs [2] or Swin Transformers [21], applying ViT blocks only within deep, low-resolution latent spaces [2, 15].
@@ 2.1 Preliminaries
where $\mathcal{R}$ represents a class of regularizers, and $\lambda$ is a balancing parameter. This formulation aims to minimise the discrepancy between the measured and predicted data while incorporating regularisation to impose prior knowledge or desired properties on the solution.
@@
Characterized by linear complexity and enhanced through hardware-aware optimisations, Mamba exhibits superior efficiency in managing long sequence modelling. This efficiency positions Mamba as a viable contender to the prevalent self-attention mechanisms found in Transformers, particularly for tasks involving the processing of high-resolution visual data.
@@ 3.2 Implementation Details and Evaluation Metrics
Our proposed models, MambaMIR and MambaMIR-GAN, were trained on one and two NVIDIA RTX 3090 GPUs (24GB each), respectively, and evaluated on a single NVIDIA RTX 3090 GPU. Both models underwent training for 100,000 gradient steps, utilizing the Adam optimizer with a learning rate of $2\times 10^{-5}$ and a batch size of 8.
@@ 3.3 Comparison Studies
TABLE/TABLE_MRI-FastMRI TABLE/TABLE_SVCT

### v3

@@ 1 Introduction
Magnetic Resonance Imaging (MRI) and X-ray Computed Tomography (CT) are two representative medical imaging tools. Magnetic resonance imaging serves as a crucial non-invasive diagnostic tool, offering high-resolution and reproducible assessments of both structural and functional information, without exposure to radiation. Fast MRI is widely utilised to produce MR images from sub-Nyquist sampled k-space measurements, aiming to speed up the inherently slow data acquisition process [18, 10, 13]. This accelerated acquisition process can effectively reduce artifacts from voluntary and involuntary physiological movements during scanning. X-ray Computed Tomography is renowned for its ability to generate high-quality and detailed images. However, it also raises significant concerns regarding potential cancer risks due to radiation exposure [30]. Sparse-view CT (SVCT), which reconstructs images from far fewer projection views, can effectively reduce the radiation dose. Moreover, in applications such as C-arm CT or dental CT, where scanning time is predominantly constrained by the slower flat-panel detector rather than the gantry’s mechanical speed, sparse-view CT emerges as a promising solution to expedite the scanning process [24]. However, reducing the number of projection views during scanning can introduce severe artefacts in reconstructed images, which can significantly undermine clinical diagnosis. Therefore, for fast MRI, SVCT, and other medical image reconstruction tasks, a key research topic and challenge is to develop effective, efficient, and reliable reconstruction models. Therefore, for fast MRI, SVCT as well as other medical image reconstruction task, a key research topic and challenge is to develop effective, efficient and reliable reconstruction model.
@@
The rapid advancement of artificial intelligence has propelled the development and widespread application of deep learning (DL)-based medical image reconstruction. Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) [3] represent two mainstream models that have achieved remarkable success in various computer vision tasks and are widely utilised in the medical imaging field. CNNs, with their hierarchical architecture and inductive biases, excel at extracting features, especially in identifying local patterns. The shared weight mechanism renders them more parameter-efficient than multilayer perceptrons (MLPs). Various CNN-based medical image reconstruction models have been proposed for different modalities, including MRI [28, 1], CT [16, 9], Ultrasound [37], and Positron Emission Tomography (PET) [6]. However, CNNs typically exhibit local sensitivity and a lack of long-range dependency, which limits their ability to contextualise global features. Vision Transformers [3], characterised by their large receptive fields and global sensitivity, often outperform CNNs in capturing extensive contextual information. Nonetheless, their significant computational demand, driven by the self-attention mechanism’s quadratic complexity with respect to sequence length, limits their practicality for medical image reconstruction, where images often have high resolution. Moreover, utilising larger patch sizes may compromise the reconstruction of detailed information [19]. Recent Transformer-based models for medical image reconstruction have sought to mitigate these limitations by: 1) Adopting a trade-off strategy that applies the self-attention mechanism within shifting windows rather than across the entire feature map [12]; 2) Constructing hybrid models that incorporate CNNs [2] or Swin Transformers [21], applying ViT blocks only within deep, low-resolution latent spaces [2, 15].
@@ 2.1 Preliminaries
where $\mathcal{R}$ represents a class of regularisers, and $\lambda$ is a balancing parameter. This formulation aims to minimise the discrepancy between the measured and predicted data while incorporating regularisation to impose prior knowledge or desired properties on the solution.
@@
Characterised by linear complexity and enhanced through hardware-aware optimisations, Mamba exhibits superior efficiency in managing long sequence modelling. This efficiency positions Mamba as a viable contender to the prevalent self-attention mechanisms found in Transformers, particularly for tasks involving the processing of high-resolution visual data.
@@ 3.2 Implementation Details and Evaluation Metrics
TABLE/TABLE_MRI-FastMRI TABLE/TABLE_SVCT
@@
Our proposed models, MambaMIR and MambaMIR-GAN, were trained on one and two NVIDIA RTX 3090 GPUs (24GB each), respectively, and evaluated on a single NVIDIA RTX 3090 GPU. Both models underwent training for 100,000 gradient steps, utilizing the Adam optimiser with a learning rate of $2\times 10^{-5}$ and a batch size of 8.